SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the February 5 Background Remediation Workshop

  1. Scrubbing issues
  2. Tom Mattison gave a status report. A total of 60 A. hours have accumulated up to now (40 in January), still a small fraction of the goal of 400 A.hours. In one occasion, 10 A.h were collected in a single day, the slope needed to reach this goal. A vacuum history measured at a pump showed that the improvments at each activation of the TSP pumps are now smaller. It is difficult to extrapolate from these data the vacuum as a function of future A.h. The lifetime history was also presented, showed that the lifetime rapidly decreases atfer ~600 mA.

    The base pressure in the LER seems to be high , around 5 nT.

    Tom also showed how to optimize the running with BABAR on the beam line. Keeping a constant level of radiation can achieve the same number of A.h with 20% less integrated dose than running a constant current. (This model does not include dose from injections). One other lesson is that it is better radiation-wise to get a given number of A.h in a longer time scale (ie patience will be required when BABAR is there).

    Although some concern was expressed about the small number of A.h accumulated so far, no action was deemed necessary since PEP-II management is already quite aware of the importance of scrubbing and will certainly do their best, given the other constraints from the machine.

  3. Report from a Beam-Abort meeting held Wednesday 3 Feb
  4. Tom Mattison reported on a Beam Abort meeting held Feb 3. A consensus was reached to keep the SLAC Analog module as part of the Abort system for the next run and to incorporate its functionnalities into a future version of the SVTRAD board. Clones of the analog modules will be made for DCH and EMC protection.

  5. Schedule of Background Remediation Sessions at the Collaboration meeting
  6. A full set of parallel sessions will be asked in common with the Background Commissioning group. Similarly, 3 miniplenary sessions will be asked and shared between the two groups according to the needs.

  7. Status report of EMC actions items
  8. Adrian McKemey reported that the PIN diodes have been determined to have the adequate sensitivity for protecting the EMC endcap. Money has been allocated to buy these and the analog module.

  9. Report on shielding issues
  10. T. Geld showed results indicating that 1.5 X0 of lead was not enough for shielding either the miniSOB or the CsI crystal close to the miniSOB. Next test will be done with 4 X0.

    John Seeman showed the proposed installation plans for the tunnel shieldings on both sides. They will consist of

    chicanes since access must be granted. The upcoming HER (DIRC SOB) side is the more advanced and should be there on May 1. On the other side, the proposal is 6 feet of concrete, supplemented by 1 cm of lead to reach the roof.

    Guy Wormser mentioned that after discussion with Georges London, the proposal is to reexamine the options

    of extra shielding in the DIRC tunnel after some data have been taken, since a significant mechanical work is certainly required to install a large structure on the magnetic shielding.

    The Drift Chamber group will look into the Monte Carlo to see whether shielding improvements on the plug would be beneficial.

  11. Monte Carlo production issues
  12. Terry Geld reported on the status of background frames production. The item on the critical path is the LER pressure profile, which should be delivered before Feb 15. Charlie Young mentioned that the actual production would take a month using all available CPUs at SLAC and LLNL, and therefore will certainly last more than a calendar month. The need of early release of small statistics for tests was expressed.

  13. IFR action items

Jim Johnson reported on the action items for the IFR. Work is on going to shield the gaps on the endcaps chambers with ¼" lead . Further segmentation of HV on layers 16 thru 18 is under study. Large increases in current draws have also been seen in correlation with temperature rise. This effect reduces the margin even further. IFR wants to ramp down during injection. Contact will be taken with Jim Olsen so IFR is included properly in the Injection hand shake. No hardware interlock wrt injection is deemed necessary for the IFR.